OpenStudy (anonymous):

How many atoms are in 15.3 cm3 of mercury (at room temperature)?

OpenStudy (jfraser):

you will need to know the density of mercury, as well as the molar mass, and the number of atoms that make up a mole

OpenStudy (anonymous):

and how do you put them together eg density * mass +atmon? or what

OpenStudy (jfraser):

set up conversion fractions to cancel units and replace them with better ones:\[(\frac{15.3cm^3 Hg}{1}) * (\frac{g Hg}{cm^3 Hg})*(\frac{molHg}{g Hg})*(\frac{atoms Hg}{mol Hg})\] plug in the numbers, and all the units except the atoms of Hg will cancel. It better be huge, because there are a lot of atoms in 15.3cm3 of Hg